Spam Corp. is financed entirely by common stock and has a beta of 1.63. The firm is expected to generate a level, perpetual stream of earnings and dividends. The stock has a price-earnings ratio of 7.80 and a cost of equity of 12.82%. The company’s stock is selling for $32. Now the firm decides to repurchase half of its shares and substitute an equal value of debt. The debt is risk-free, with an interest rate of 5%. The company is exempt from corporate income taxes. Assume MM are correct.

a. Calculate the cost of equity after the refinancing. (Enter your answer as a percent rounded to 2 decimal places.)

Cost of equity             %  

b. Calculate the overall cost of capital (WACC) after the refinancing. (Enter your answer as a percent rounded to 2 decimal places.)

Cost of capital             %

c. Calculate the price-earnings ratio after the refinancing. (Round your answer to 2 decimal places.)

Price-earnings ratio            

d. Calculate the stock price after the refinancing. (Round your answer to the nearest whole number.)

Stock price            $

e. Calculate the stock’s beta after the refinancing. (Round your answer to 1 decimal place.)

Stock's beta            
Answer #1

a. Given a 12.5% cost of equity before debt, we find the expected cost on equity:

= 0.5 x K(e) + 0.5 x 5% = 12.82%

= K(e) = 20.64%

b. The overall cost of capital will remain at 12.82%.

c. Maintaining the perpetual stream of earnings and dividends, the inverse of the PE ratio must now be E/P = 20.64%, which implies a P/E ratio of = 4.84.

d. Assuming MM is correct, the stock price remains at $32.

e. Since the debt is risk free, its beta is zero; the beta of the stock is found then by:

= 0.5 X equity beta + 0.5 X 0 = 1.0

= Equity beta = 2.0